A ball is thrown from 4 feet off the ground with a vertical velocity of 30 feet per second.
Is the ball at its maximum height after 1 second? In two or more complete sentences, explain how you know whether or not the ball is at its maximum height.

Answer by Boreal(15235)

-16t^2+30t+4
after 1 second-16+30+4=18 feet
maximum height is when t=-b/2a, or 30/32 (15/16)
-16(225/256)-450/16+4
=(-3600+7200+1024)/256=4624/256=18.0625'
This can be described in words. We would know when the time (15/16) is not the same as 1 second.